# Prompt-OT: An Optimal Transport Regularization Paradigm for Knowledge Preservation in Vision-Language Model Adaptation

### v1

@@ Abstract
Vision-language models (VLMs) such as CLIP demonstrate strong performance but struggle when adapted to downstream tasks. Prompt learning has emerged as an efficient and effective strategy to adapt VLMs while preserving their pre-trained knowledge. However, existing methods still lead to overfitting and degrade zero-shot generalization. To address this challenge, we propose an optimal transport (OT)-guided prompt learning framework that mitigates forgetting by preserving the structural consistency of feature distributions between pre-trained and fine-tuned models. Unlike conventional point-wise constraints, OT naturally captures cross-instance relationships and expands the feasible parameter space for prompt tuning, allowing a better trade-off between adaptation and generalization. Our approach enforces joint constraints on both vision and text representations, ensuring a holistic feature alignment. Extensive experiments on benchmark datasets demonstrate that our simple yet effective method can outperform existing prompt learning strategies in base-to-novel generalization, cross-dataset evaluation, and domain generalization without additional augmentation or ensemble techniques. The code is available at https://github.com/ChongQingNoSubway/Prompt-OT
@@ 1 Introduction
Foundational vision-language models (VLMs), such as CLIP [19] and ALIGN [8], trained on large-scale datasets to align image-text pairs have demonstrated exceptional generalization capabilities across a variety of downstream tasks [11, 4, 15, 31]. However, recent studies have shown that VLMs face challenges in maintaining generalization when adapted to downstream tasks via e.g. fine-tuning [10], particularly for tasks with limited samples such as few-shot learning.
@@
However, we argue that this may not be an optimal way to impose the aforementioned feature consistency due to the following reasons: (i) point-wise constraints may fail to capture cross-instance relationships, limiting their ability to account for the geometry of the underlying manifold and effectively model both within-class and cross-class relations. (ii) Point-wise constraints might be too restrictive111 A similar conclusion is also drawn in a well-known work [12] for related tasks. We have more discussion for this method in Sec. 2, as different features and samples have different importance. This may cause the features to learn from prompted VLM to be shrunk to a narrow searching space. This limitation can hinder both adaptation and generalizability.
@@
To address these challenges, we propose an optimal transport (OT)-guided VLMs prompt tuning method, which leverages OT to mitigate knowledge forgetting between the embedding features of the zero-shot and the finetuned VLMs. The advantages of OT-guided prompts are multi-fold. First, OT naturally captures cross-instance relationships, since the transport map establishes an interactive coupling among instances by redistributing mass based on their mutual relationships. Second, we prove that OT expands the feasible domain for learnable prompt tokens. A broader feasible domain introduces more potential local minima, provides greater flexibility for the optimizer to explore, and allows easier hyperparameter adjustment, increasing the likelihood of achieving an optimal balance between adaptation and generalization. The conceptual illustration is shown in Fig. 1. Third, rather than independently constraining text and vision representations and merely computing a weighted sum of their losses, we incorporate the fundamental concept of an instance from CLIP, which inherently integrates both vision and text modalities. Consequently, our design enforces constraints on both modalities simultaneously for each instance. It is also worth noting that, despite its simple yet effective design, our method outperforms several recent competitive approaches without relying on additional components such as ensemble methods or data augmentation techniques. In summary, our work makes the following key contributions:
  • We propose an OT-based constraint scheme to mitigate knowledge forgetting, addressing the core limitations of previous constraint-based prompt learning methods.
  • Inspired by the fundamental concept of an instance in CLIP, we impose constraints on both modalities simultaneously for each instance.
  • Empirical evaluations across three popular benchmarks (i.e., Base-To-Novel Generalization, Cross-Dataset Evaluation, and Domain Generalization), along with ablation studies, demonstrate the effectiveness and superiority of our approach over recent alternative methods, such as PromptSRC [10], TCP [26], and QMaPLe [7].
@@ 2 Related Work
Adaption to Downstream tasks. Two popular ways to adapt a model for downstream tasks are full fine-tuning and linear probing. However, full fine-tuning can weaken generalization, while linear probing often yields poor performance, leading to both being problematic for tasks needing accuracy and generalization, like open-vocabulary image recognition and object detection [9]. To address this, researchers add extra learnable components without altering the original pre-trained weights, attempting to preserve both performance and generalization.[29]. One representative approach is known as prompt-based methods [29, 28, 30, 26]. In these methods, the pre-trained model is adapted by incorporating a small number of learnable prompt tokens into the original input. For example, CoOp [29], CoCoOp [28], ProdGrad [30], TCP [26] introduced a set of learnable continuous vectors into the text input. Although there are some attempts to only include prompt tokens on the image branch, more recent works prefer to perform prompt learning on both branches jointly, such as MaPLe [9] and PromptSRC [10]. Instead, adapter-based methods often add some extra layers to refine features generated by CLIP [27, 6, 25].
Consistency-based Learning. Instead of focusing on the design of the architecture, there are methods to enforce consistency between the trainable and pre-trained models to reduce overfitting and forgetting on downstream task fine-tuning. ProGrad [30] aligns gradient directions with the pre-trained model, while PromptSRC [10] imposes consistency on both embeddings and logits. A similar idea in [22] employs adapters, data augmentations, and LLM knowledge. Though PromptSRC is closest to our method, it only applies pairwise constraints that match average embedding distributions. We argue this is too restrictive and overlooks structural information, as discussed in Sec. 3.3. A related work [12] in continual learning attempts to solve severe restriction issues through a Fisher Information-based constraint, while it requires a substantial approximation during computing and needs the corresponding weights on the frozen model, which makes it infeasible to our case as the prompt tokens are only adding to the adapted model.
@@
Augmentation for VLMs. Several recent studies introduce extra augmentation to boost VLMs adaptation. For example, [10, 26] use more hand-crafted text templates beyond “a photo of a [category]” or design their own prompts, while [22, 32] leverage LLMs for diverse semantic prompts. Others [5] rely on Stable Diffusion [21] to generate multi-view images for better test-time tuning. These methods show that high-quality augmentation can significantly improve performance on downstream tasks. However, not everyone can access such data-augmentation resources. Thus, our method is designed to work without any extra augmentation, yet still delivers competitive results.
@@ 3 Method
Figure 2: The overview of our proposed framework. Only the prompt tokens are trainable, and the rest of the weights in both the zero-shot encoders and the adapted encoders are frozen. Despite the cross-entropy $\mathcal{L}_{ce}$ adopted, we also adopt the proposed joint optimal transport loss $\mathcal{L}_{\mathrm{jot}}$ between joint zero-short representation and adapted representation to constrain the model.
@@ 3.1 Preliminaries
Vision Encoding. An image $I$ is first divided into $M$ patches and then projected to latent space, which is denoted as $\bm{V}^{0}=\{\bm{v}_{1}^{0},\cdots,\bm{v}_{M}^{0}\}$. Afterward, a learnable class token $\bm{v}_{cls}^{0}$ is concatenated as the global representation, denoted as $[\bm{V}^{0},\bm{v}_{cls}^{0}]$. These tokens subsequently are fed into a sequence of transformer blocks:
$[\bm{V}^{l},\bm{v}_{cls}^{l}]$ $=\texttt{Trans}^{l}_{\texttt{Vision}}([\bm{V}^{l-1},\bm{v}_{cls}^{l-1}]),$ (1)
where $l=1,\cdots,L_{I}$ and $L_{I}$ denote the number of vision transformer blocks in the image encoder of CLIP. Afterward, the final image representation in V-L space is obtained through an additional linear projection:
$\bm{h}=f_{\texttt{VisionProj}}(\bm{v}_{cls}^{L_{I}}).$ (2)
Text Encoding. Each text prompt is tokenized to $N$ tokens, denoted as $\bm{T}^{0}=\{\bm{t}_{1}^{0},\cdots,\bm{t}_{N}^{0}\}$. Likewise, these tokens then feed to text transformer blocks,
$\bm{T}^{l}$ $=\texttt{Trans}^{l}_{\texttt{Text}}(\bm{T}^{l-1}),$ (3)
where $l=1,\cdots,L_{T}$ and $L_{T}$ denote the number of transformer blocks in the text encoder of CLIP. Similarly, the final representation is obtained by the last token through a linear projection:
$\bm{g}=f_{\texttt{TextProj}}(\bm{t}_{N}^{L_{T}}).$ (4)
@@
$p(\bar{y}|I)=\frac{\text{exp}(sim(\bm{h},\bm{g}_{y})/\tau)}{\sum_{k=1}^{C}\text{exp}(sim(\bm{h},\bm{g}_{i}/\tau)},$ (5)
@@
Prompt Learning. Now, we will delve into how to perform prompt learning, which often appends some learnable tokens to text or vision branches as some transformer layers. Formally, we denote $\mathcal{S}_{\texttt{vision}}$ and $\mathcal{S}_{\texttt{text}}$ as the layers to integrate learnable prompt tokens for vision encoder and text encoder, respectively. We can denote the $K_{I}$ vision prompt tokens and $K_{T}$ text prompt tokens at $(l+1)$-th block as $\bm{P}^{l}=\{\bm{p}^{l}_{1},\cdots,\bm{p}^{l}_{K_{I}}\}$ and $\bm{Q}^{l}=\{\bm{q}^{l}_{1},\cdots,\bm{q}^{l}_{K_{T}}\}$, respectively. The prompting for vision encoder in $l$-th layer ($l\in\mathcal{S}_{\texttt{vision}}$) can be then presented as:
[¯$[~{}\underline{\hskip 8.5359pt}~{},\bm{V}^{l},\bm{v}_{cls}^{l}]$ $=\texttt{Trans}^{l}_{\texttt{Vision}}([\bm{P}^{l-1},\bm{V}^{l-1},\bm{v}_{cls}^{l-1}]),$ (6)
@@
[¯$[~{}\underline{\hskip 8.5359pt}~{},\bm{T}^{l}]$ $=\texttt{Trans}^{l}_{\texttt{Text}}([\bm{Q}^{l-1},\bm{T}^{l-1}]).$ (7)
@@ 3.2 Vision-Text Joint Optimal Transport
In this section, we will delve into the design of our framework. During training, this framework involves a CLIP with learnable prompts, i.e., the adapted model and a frozen pre-trained CLIP, while during inference, we only need the adapted CLIP. The overview of this framework is shown in Fig. 2.
A key principle in CLIP is that an instance has representations from both the text domain and the image domain. Therefore, we should constrain them simultaneously for each instance. Suppose we have a dataset $\mathcal{D}=\{(I_{i},y_{i})\}_{i=1}^{n}$, where $I_{i},y_{i}$, and $n$ denote the $i$th input image, its corresponding label, and number of samples, respectively. For each instance $i\in[n]$, the vision representation is obtained by the vision encoder as $\bm{h}^{i}$. Since the real text caption for the image is often unknown, we use the standard template with its category as the proxy, and accordingly, we can obtain the text representation as $\bm{g}^{i}$. For simplicity, we concatenate both representations as the joint representation of an instance $i$, denoted as $\bm{x}_{i}=\texttt{concat}(\bm{h}^{i},\bm{g}^{i})$. Now, we have the distribution of joint representation of the dataset $\mathcal{D}$ as $\mathcal{X}=\{\bm{x}^{i}\}_{i=1}^{n}$ by the adapted model. We also need to obtain the joint representation distribution using an additional CLIP model, i.e., the CLIP model with its original pre-trained weights, while keeping all parameters frozen. We denote it as $\bm{X}_{\texttt{zs}}=\{\bm{x}_{\texttt{zs}}^{i}\}_{i=1}^{n}$. It is worth mentioning that $\bm{x}^{i}$ and $\bm{x}_{\texttt{zs}}^{i}$ are obtained from the same input instance $(I_{i},y_{i})$, while with different CLIP models (i.e. one keep updating and one frozen).
Our proposed regularization term is presented as: {definition} Given $\bm{X}$ and $\bm{X}_{\texttt{zs}}$, the proposed constraint lies in Optimal Transport [23] is defined as:
$\mathcal{L}_{\mathrm{jot}}(\bm{X},\bm{X}_{\texttt{zs}})=\min_{\bm{\gamma}\in\mathcal{U}(\bm{a},\bm{b})}\sum_{i=1}^{n}\sum_{j=1}^{n}\bm{\gamma}_{ij}\,c(\bm{x}^{i},\bm{x}_{\texttt{zs}}^{j}),$ (8)
where $c(\bm{x}^{i},\bm{x}_{\texttt{zs}}^{j})$ denotes a cost function (e.g. linear and quadratic cost) between $\bm{x}^{i}$ and $\bm{x}_{\texttt{zs}}^{j}$ $\bm{\gamma}$ denotes a transport map, $\mathcal{U}(\bm{a},\bm{b})$ denotes set of all valid transport maps,
$\mathcal{U}(\bm{a},\bm{b})$ =$=\left\{\bm{\gamma}\in\mathbb{R}^{n\times n}_{+}\ \middle|\ \bm{\gamma}\mathbf{1}_{n}=\bm{a},\ \bm{\gamma}^{\top}\mathbf{1}_{n}=\bm{b}\right\},$ (9)

### v2

@@ Abstract
Vision-language models (VLMs) such as CLIP demonstrate strong performance but struggle when adapted to downstream tasks. Prompt learning has emerged as an efficient and effective strategy to adapt VLMs while preserving their pre-trained knowledge. However, existing methods still lead to overfitting and degrade zero-shot generalization. To address this challenge, we propose an optimal transport (OT)-guided prompt learning framework that mitigates forgetting by preserving the structural consistency of feature distributions between pre-trained and fine-tuned models. Unlike conventional point-wise constraints, OT naturally captures cross-instance relationships and expands the feasible parameter space for prompt tuning, allowing a better trade-off between adaptation and generalization. Our approach enforces joint constraints on both vision and text representations, ensuring a holistic feature alignment. Extensive experiments on benchmark datasets demonstrate that our simple yet effective method outperforms existing prompt learning strategies in base-to-novel generalization, cross-dataset evaluation, and domain generalization, without requiring additional augmentation or ensemble techniques. The code is available at https://github.com/ChongQingNoSubway/Prompt-OT
@@ 1 Introduction
Foundational vision-language models (VLMs), such as CLIP [30] and ALIGN [15], trained on large-scale datasets to align image-text pairs, have demonstrated exceptional generalization capabilities across a variety of downstream tasks [18, 8, 23]. However, recent studies have shown that VLMs face challenges in maintaining generalization when adapted to downstream tasks via e.g. fine-tuning [17], particularly for tasks with limited samples such as few-shot learning.
@@
Despite this, we argue that this may not be an optimal way to impose the aforementioned feature consistency due to the following reasons: (i) Point-wise constraints may fail to capture cross-instance relationships, limiting their ability to account for the geometry of the underlying manifold and effectively model both within-class and cross-class relations. (ii) Point-wise constraints might be too restrictive [^1] A similar conclusion is also drawn in a well-known work [19] for related tasks. We have more discussion for this method in Sec. 2., as different features and samples have different importance. This may cause the features to learn from prompted VLM to be shrunk to a narrow searching space. This limitation can hinder both adaptation and generalizability.
@@
To address these challenges, we propose an optimal transport (OT)-guided VLMs prompt tuning method, which leverages OT to mitigate knowledge forgetting between the embedding features of the zero-shot and the finetuned VLMs. The advantages of OT-guided prompts are multi-fold. First, OT naturally captures cross-instance relationships, since the transport map establishes an interactive coupling among instances by redistributing mass based on their mutual relationships. Second, we prove that OT expands the feasible domain for learnable prompt tokens. A broader feasible domain introduces more potential local minima, provides greater flexibility for the optimizer to explore, and allows easier hyperparameter adjustment, increasing the likelihood of achieving an optimal balance between adaptation and generalization. The conceptual illustration is shown in Fig. 1. Third, existing methods often constrain vision and text representations independently and then combine them with a weighted sum of losses. However, the vision encoder outputs features of size $B\times d$ (where $B$ and $d$ denote the batch size and the latent dimension, respectively), while the text encoder outputs features of size $C\times d$ (where $C$ denotes the number of classes). This mismatch can lead to imbalanced constraint strength across modalities. To address this, we build on the fundamental concept of an instance in CLIP, which inherently couples both vision and text. Our design therefore enforces constraints on vision and text simultaneously at the instance level, ensuring more balanced and structurally consistent multimodal representations. Consequently, our design enforces constraints on both modalities simultaneously for each instance. It is also worth noting that, despite its simple yet effective design, our method outperforms several recent competitive approaches without relying on additional components such as ensemble methods or data augmentation techniques. In summary, our work makes the following key contributions:
  • We propose an OT-based constraint scheme to mitigate knowledge forgetting, addressing the core limitations of previous constraint-based prompt learning methods. Beyond the conceptual design, we provide a theoretical justification showing that OT enlarges the feasible parameter space of learnable prompts, thereby enabling a more flexible and principled trade-off between adaptation and generalization.
  • We propose to impose the OT constraint jointly on vision and text features of each instance. This balances the constraint strength across modalities and preserves structural consistency in multimodal representations, addressing the limitations of treating modalities independently.
  • Empirical analysis (Table 4) highlights that Prompt-OT consistently outperforms all point-wise constraint methods with statistical significance. Across three popular benchmarks (Base-to-Novel Generalization, Cross-Dataset Evaluation, and Domain Generalization), our method achieves competitive or superior performance compared to recent approaches, including PLOT [2], PromptSRC [17], TCP [41], and QMaPLe [11], without relying on prompt ensembles or extra augmentation.
@@ 2 Related Work
Adaptation to Downstream tasks. Two popular ways to adapt a model for downstream tasks are full fine-tuning and linear probing. However, full fine-tuning can weaken generalization, while linear probing often yields poor performance, leading to both being problematic for tasks needing accuracy and generalization, like open-vocabulary image recognition and object detection [16]. To address this, researchers add extra learnable components without altering the original pre-trained weights, attempting to preserve both performance and generalization [44]. One representative approach is known as prompt-based methods [44, 43, 45, 41]. In these methods, the pre-trained model is adapted by incorporating a small number of learnable prompt tokens into the original input. For example, CoOp [44], CoCoOp [43], ProdGrad [45], TCP [41] introduced a set of learnable continuous vectors into the text input. Although there are some attempts to only include prompt tokens on the image branch, more recent works prefer to perform prompt learning on both branches jointly, such as MaPLe [16] and PromptSRC [17]. Instead, adapter-based methods often add some extra layers to refine features generated by CLIP [42, 10, 40].
Consistency-based Learning. Instead of focusing on the design of the architecture, there are methods to enforce consistency between the trainable and pre-trained models, thereby reducing overfitting and forgetting during downstream task fine-tuning. ProGrad [45] aligns gradient directions with the pre-trained model, while PromptSRC [17] imposes consistency on both embeddings and logits. A similar idea in [34] employs adapters, data augmentations, and LLM knowledge. Though PromptSRC is closest to our method, it only applies pairwise constraints that match average embedding distributions. We argue this is too restrictive and overlooks structural information, as discussed in Sec. 3.3. A related work [19] in continual learning attempts to solve severe restriction issues through a Fisher Information-based constraint, while it requires a substantial approximation during computing and needs the corresponding weights on the frozen model, which makes it infeasible for our case, as the prompt tokens are only added to the adapted model. Additionally, previous methods typically involve terms that constrain each modality independently. However, during CLIP training, the feature spaces of the two modalities are inherently imbalanced: the image encoder produces a $B\times d$ feature map, while the text encoder outputs a fixed $C\times d$ representation. This discrepancy leads to uneven constraint strength across modalities. To address this, we propose imposing constraints directly on the joint representation of each instance, ensuring better alignment. A detailed analysis is provided in Sec. 5 and Table 5.
@@
Augmentation for VLMs. Several recent studies introduce extra augmentation to boost VLMs’ adaptation. For example, [17, 41] use more hand-crafted text templates beyond “a photo of a [category]” or design their prompts, while [34, 46] leverage LLMs for diverse semantic prompts. Others [9] rely on Stable Diffusion [33] to generate multi-view images for better test-time tuning. These methods show that high-quality augmentation can significantly improve performance on downstream tasks. However, not everyone can access such data-augmentation resources. Thus, our method is designed to work without any extra augmentation, yet still delivers competitive results.
@@ 3.1 Preliminaries
Vision Encoding. An image $I$ is first divided into $M$ patches and then projected to latent space, which is denoted as $\bm{V}^{0}=\{\bm{v}_{1}^{0},\cdots,\bm{v}_{M}^{0}\}$. Afterward, a learnable class token $\bm{v}_{cls}^{0}$ is concatenated as the global representation, denoted as $\bm{V}^{0},\bm{v}_{cls}^{0}]$. These tokens are subsequently fed into a sequence of transformer blocks: $\bm{V}^{l},\bm{v}_{cls}^{l}]=\texttt{Trans}^{l}_{\texttt{Vision}}([\bm{V}^{l-1},\bm{v}_{cls}^{l-1}]),$ Where $l=1,\cdots,L_{I}$ and $L_{I}$ denote the number of vision transformer blocks in the image encoder of CLIP. Afterward, the final image representation in V-L space is obtained through an additional linear projection: $\bm{h}=f_{\texttt{VisionProj}}(\bm{v}_{cls}^{L_{I}}).$
Text Encoding. Each text prompt is tokenized to $N$ tokens, denoted as $\bm{T}^{0}=\{\bm{t}_{1}^{0},\cdots,\bm{t}_{N}^{0}\}$. Likewise, these tokens are then fed to text transformer blocks as $\bm{T}^{l}=\texttt{Trans}^{l}_{\texttt{Text}}(\bm{T}^{l-1}),$ Where $l=1,\cdots,L_{T}$ and $L_{T}$ denote the number of transformer blocks in the text encoder of CLIP. Similarly, the final representation is obtained by the last token through a linear projection as $\bm{g}=f_{\texttt{TextProj}}(\bm{t}_{N}^{L_{T}}).$
@@
$p(\bar{y}|I)=\frac{\text{exp}(sim(\bm{h},\bm{g}_{y})/\tau)}{\sum_{k=1}^{C}\text{exp}(sim(\bm{h},\bm{g}_{k}/\tau))},$ (1)
@@
Prompt Learning. Now, we will delve into how to perform prompt learning, which often appends some learnable tokens to text or vision branches as part of transformer layers. Formally, we denote $\mathcal{S}_{\texttt{vision}}$ and $\mathcal{S}_{\texttt{text}}$ as the layers to integrate learnable prompt tokens for the vision encoder and text encoder, respectively. We can denote the $K_{I}$ vision prompt tokens and $K_{T}$ text prompt tokens at $(l+1)$-th block as $\bm{P}^{l}=\{\bm{p}^{l}_{1},\cdots,\bm{p}^{l}_{K_{I}}\}$ and $\bm{Q}^{l}=\{\bm{q}^{l}_{1},\cdots,\bm{q}^{l}_{K_{T}}\}$, respectively. The prompting for vision encoder in $l$-th layer ($l\in\mathcal{S}_{\texttt{vision}}$) can be then presented as:
$[~\underline{\hskip 8.5359pt}~,\bm{V}^{l},\bm{v}_{cls}^{l}]$ $=\texttt{Trans}^{l}_{\texttt{Vision}}([\bm{P}^{l-1},\bm{V}^{l-1},\bm{v}_{cls}^{l-1}]),$ (2)
@@
$[~\underline{\hskip 8.5359pt}~,\bm{T}^{l}]$ $=\texttt{Trans}^{l}_{\texttt{Text}}([\bm{Q}^{l-1},\bm{T}^{l-1}]).$ (3)
@@ 3.2 Vision-Text Joint Optimal Transport
In this section, we will delve into the design of our framework. During training, this framework involves a CLIP with learnable prompts, i.e., the adapted model and a frozen pre-trained CLIP, while during inference, we only need the adapted CLIP.
As discussed in Sec. 2, constraining each modality separately can result in uneven constraint strength due to inherent imbalances. To address this, we emphasize a key principle of CLIP: each instance is jointly represented in both the text and image domains. Consequently, the two modalities should be constrained simultaneously at the instance level. Suppose we have a dataset $\mathcal{D}=\{(I_{i},y_{i})\}_{i=1}^{n}$, where $I_{i},y_{i}$, and $n$ denote the $i$th input image, its corresponding label, and the number of samples, respectively. For each instance $i\in[n]$, the vision representation is obtained by the vision encoder as $\bm{h}^{i}$. Since the real text caption for the image is often unknown, we use the standard template with its category as the proxy, and accordingly, we can obtain the text representation as $\bm{g}^{i}$. For simplicity, we concatenate both representations as the joint representation of an instance $i$, denoted as $\bm{x}_{i}=\texttt{concat}(\bm{h}^{i},\bm{g}^{i})$. Now, we have the distribution of joint representation of the dataset $\mathcal{D}$ as $\mathcal{X}=\{\bm{x}^{i}\}_{i=1}^{n}$ by the adapted model. We also need to obtain the joint representation distribution using an additional CLIP model, i.e., the CLIP model with its original pre-trained weights, while keeping all parameters frozen. We denote it as $\bm{X}_{\texttt{zs}}=\{\bm{x}_{\texttt{zs}}^{i}\}_{i=1}^{n}$. It is worth mentioning that $\bm{x}^{i}$ and $\bm{x}_{\texttt{zs}}^{i}$ are obtained from the same input instance $(I_{i},y_{i})$, while with different CLIP models (i.e., one keeps updating and one is frozen). Our proposed regularization term is presented as follows: Given $\bm{X}$ and $\bm{X}_{\texttt{zs}}$, the proposed constraint lies in Optimal Transport [36] is defined as:
$\vskip-7.22743pt\mathcal{L}_{\mathrm{jot}}(\bm{X},\bm{X}_{\texttt{zs}})=\min_{\bm{\gamma}\in\mathcal{U}(\bm{a},\bm{b})}\sum_{i=1}^{n}\sum_{j=1}^{n}\bm{\gamma}_{ij}\,c(\bm{x}^{i},\bm{x}_{\texttt{zs}}^{j}),\vskip-7.22743pt$ (4)
where $c(\bm{x}^{i},\bm{x}_{\texttt{zs}}^{j})$ denotes a cost function (e.g. linear and quadratic cost) between $\bm{x}^{i}$ and $\bm{x}_{\texttt{zs}}^{j}$ . $\bm{\gamma}$ denotes a transport map, $\mathcal{U}(\bm{a},\bm{b})$ denotes set of all valid transport maps,
$\mathcal{U}(\bm{a},\bm{b})=\left\{\bm{\gamma}\in\mathbb{R}^{n\times n}_{+}\ \middle|\ \bm{\gamma}\mathbf{1}_{n}=\bm{a},\ \bm{\gamma}^{\top}\mathbf{1}_{n}=\bm{b}\right\},$ (5)